OpenStudy (jewlzme17):

Solve using two different methods. Explain which method you found to be efficient. A. 3x-9y= 3 B. 7x-3y= 20 C. y=1/2x-6 6x-3y= -24 5x+3y=16 2x+6y= 19

jhonyy9 (jhonyy9):

B elimination method add term by term 7x-3y=20 5x+3y=16 ----------- 12x 0 =36 x=36/12 =3 x=3 hope you can continue it

jhonyy9 (jhonyy9):

A. multiplie first equation by -2 and will get -6x+18y=-6 6x-3y=-24 ------------- assum it 0 15y=-30 y=-30/15=-2 y=-2 hope so much that you will can continue it

jhonyy9 (jhonyy9):

C. substitute the value of y from the first equation in place of y in second equation good luck bye hope that i was in your help

OpenStudy (jewlzme17):

Thank you for your help it helped me a little bit more

OpenStudy (turingtest):

which one would you like to do?

OpenStudy (jewlzme17):

I have to do all three of them...

OpenStudy (turingtest):

well I can't show you how to do all three both ways, but I think I have time to walk you through one hopefully that will help you do the others

OpenStudy (turingtest):

so pick one...

OpenStudy (jewlzme17):

ok thank you

OpenStudy (jewlzme17):

C

OpenStudy (turingtest):

multiply the first equation by 2 what do you get ?

OpenStudy (jewlzme17):

idk i am so confused.........?

OpenStudy (turingtest):

sorry, I meant to say 4 y=1/2x-6 2x+6y=19 multiply the first equation by 4 and you get 4y=2x-24 2x+6y=19 now get the first equation arranged so that x and y are on the left side of the equation, and the constant -24 is on the right

OpenStudy (jewlzme17):

Oh okay